Measures of liquidity, Solvency, and Profitability

The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $ 63 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $1,022,050 $862,950
Net income 233,600 176,800
Total $1,255,650 $1,039,750
Dividends:
On preferred stock $7,700 $7,700
On common stock 10,000 10,000
Total dividends $17,700 $17,700
Retained earnings, December 31 $1,237,950 $1,022,050


Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $1,470,220 $1,354,620
Cost of goods sold 571,590 525,860
Gross profit $898,630 $828,760
Selling expenses $291,610 $366,610
Administrative expenses 248,400 215,310
Total operating expenses $540,010 $581,920
Income from operations $358,620 $246,840
Other revenue 18,880 15,760
$377,500 $262,600
Other expense (interest) 112,000 61,600
Income before income tax $265,500 $201,000
Income tax expense 31,900 24,200
Net income $233,600 $176,800


Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   20Y2    20Y1
Assets
Current assets
Cash $345,040 $199,520
Marketable securities 522,220 330,630
Accounts receivable (net) 284,700 270,100
Inventories 219,000 175,200
Prepaid expenses 65,275 39,900
Total current assets $1,436,235 $1,015,350
Long-term investments 610,460 75,113
Property, plant, and equipment (net) 1,540,000 1,386,000
Total assets $3,586,695 $2,476,463
Liabilities
Current liabilities $478,745 $214,413
Long-term liabilities:
Mortgage note payable, 8% $630,000 $0
Bonds payable, 8% 770,000 770,000
Total long-term liabilities $1,400,000 $770,000
Total liabilities $1,878,745 $984,413
Stockholders' Equity
Preferred $0.70 stock, $20 par $220,000 $220,000
Common stock, $10 par 250,000 250,000
Retained earnings 1,237,950 1,022,050
Total stockholders' equity $1,707,950 $1,492,050
Total liabilities and stockholders' equity $3,586,695 $2,476,463

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

1. Working capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Number of days' sales in receivables days
6. Inventory turnover
7. Number of days' sales in inventory days
8. Ratio of fixed assets to long-term liabilities
9. Ratio of liabilities to stockholders' equity
10. Times interest earned
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield %
Answer #1

1) Working capital= Current assets-Current liabilities

= $1436235-478745

= $957490

2) Current ratio= Current assets/Current liabilities

= $1436235/478745

= 3

3) Quick assets= Cash+Marketable securities+Accounts receivable

= $345040+522220+284700= $1151960

Quick ratio= Quick assets/Current liabilities

= $1151960/478745= 2.4

4) Average account receivables= ($284700+270100)/2 = $277400

Accounts receivable turnover= Sales/Average account receivables

= $1470220/277400= 5.3 times

Note:- As per HOMEWORKLIB RULES, we can only solve first four parts in case of multiple parts. So please post all the other parts in the group of four four parts separately